# Tracking the variation of entanglement Rényi negativity : A quantum Monte Carlo study

### v2

@@ Abstract
Introduction.- The increasing connection between condensed matter physics and quantum information science have significantly deepened our understanding of quantum many-body behaviors. Beyond the conventional Landau-Ginzburg-Wilson framework of symmetry breaking and phase transitions Sachdev (1999); Lifshitz and Pitaevskii (2013), quantum entanglement and other intricate microcosmic degrees of freedom have become central in explaining various exotic emergent phenomena Amico et al. (2008a); Laflorencie (2016); Zeng et al. (2019). In a bipartite composite system, the entanglement of a pure ground state can be quantified by entanglement entropy (EE), which captures universal information of the system including critical behaviors Kallin et al. (2014); Helmes and Wessel (2014); Zhao et al. (2022a, b), continuous symmetry breaking Metlitski and Grover (2011); D’Emidio (2020); Deng et al. (2023); Wang et al. (2024a), the underlying conformal field theory (CFT) Calabrese and Lefevre (2008); Fradkin and Moore (2006); Casini and Huerta (2007); Ji and Wen (2019); Tang and Zhu (2020) and topological order Kitaev and Preskill (2006); Levin and Wen (2006); Isakov et al. (2011); Nussinov and Ortiz (2009a, b). However, EE fails as a measure of entanglement in mixed states, which frequently arise in many-body physics, such as in finite-temperature Gibbs states, systems with disconnected partitions, and open quantum systems Lifshitz and Pitaevskii (2013); Laflorencie (2016); Yan et al. (2018). Quantifying mixed-state entanglement remains difficult both analytically and numerically, as computing most existing measures requires complex optimization over the full Hilbert space Huang (2014); Amico et al. (2008b); Horodecki et al. (2009); Eisert and Plenio (1999).
@@
In recent years, some numerical approaches have been developed to calculate RN or the moments Alba (2013); Shim et al. (2018); Chung et al. (2014a); Wu et al. (2020a); Wang and Xu (2023). However, efficient algorithms for exploring how RN varies with system parameters in large-scale and high-dimensional settings remain scarce, as existing methods typically compute RN at only a single parameter point per run, making it computationally expensive to obtain a full RN curve. Particularly for large-scale interacting spin/boson systems, there are few quantum Monte Carlo (QMC) works Chung et al. (2014b); Wu et al. (2020b), mainly because the incremental trick commonly used in calculating Rényi EE Hastings et al. (2010); Humeniuk and Roscilde (2012); D’Emidio (2020) becomes much more complex and difficult in the case of RN, since the lowest order of nontrivial RN is third and the boundary condition of imaginary time is twisted. Therefore, developing an efficient QMC algorithm that is both easy to implement and capable of accurately computing RN is a crucial yet challenging task. Furthermore, the absence of such algorithms has left many important questions about mixed-state entanglement unexplored.
In this letter, we fill the gap by presenting an efficient QMC algorithm within the reweight-annealing frame Ding et al. (2024); Wang et al. (2024b); Jiang et al. (2024); Ding et al. (2025). The named annealing scheme, a specific incremental approach in the algorithm, allows for high precision while maintaining ease of implementation, as it avoids modifications to the entangled regions and boundary condition of imaginary time during simulations. Moreover, the algorithm can naturally generate a continuous RN curve across parameter space and is inherently parallelizable. It enables us to track the variation of RN with respect to a relevant parameter, which is important for understanding many-body entanglement in different phases and at critical points. We demonstrate its capacity in (i) the tripartite 1D antiferromagnet dimerized Heisenberg model (ADHM) at zero temperature, (ii) the bipartite 1D and 2D transverse field Ising models (TFIM) by varying the external field strength at zero temperature, and (iii) the bipartite 2D TFIM by adjusting temperature with weak external fields. Not only the information of the underlying CFTs are achieved, but the roles of entanglement in quantum and thermal phase transitions are revealed in this work.
@@
Figure 1: The horizontal axis represents spatial spins (depicted as circles), while the vertical axis corresponds to imaginary time. Spins of the same color are connected (or glued) along the imaginary time direction by straight lines of the same color. (a) Manifold for $Z_{3}$: The three replicas are sequentially connected along the imaginary time direction for the system $A\cup B$. The environment, denoted by $\overline{A\cup B}$, is traced out by imposing periodic boundary conditions on each replica. (b) manifold for $Z_{n}^{T_{B}}$: The connections among three replicas in subsystem $A$ (the blue part) and the environment $\overline{A\cup B}$ (the yellow part) remain identical to that in the left manifold. However, in subsystem $B$, the connections are not in sequence due to the operation $T_{B}$.
@@
The ratio $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ in Eq. (2) can be further interpreted as a reweighting operation Ferrenberg and Swendsen (1988, 1989); Troyer et al. (2004). If express $Z_{n}^{T_{B}}(\lambda)=\sum_{i}W_{i}(\lambda)$, where $\{W_{i}(\lambda)\}$ are the weights, then the ratio can be viewed as the average of weight ratios, since $Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})=\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$, where $W(\lambda)\in\{W_{i}(\lambda)\}$ and the expectation is computed with respect to samples from $Z_{n}^{T_{B}}(\lambda^{\prime\prime})$. The specific form of the estimator $\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$ depends on the QMC method we use, which transforms the quantum degrees of freedom in $Z_{n}^{T_{B}}(\lambda^{\prime})$ to some classical ones. For illustration, we use the stochastic series expansion (SSE) method in this letter Sandvik (1998, 2003); Yan et al. (2019); Yan (2022), and the corresponding estimators can be found in the supplemental materials (SM). Similarly, we can apply the same reasoning to $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$. It is important to emphasize that the view of reweighting is crucial, as it allows us to estimate the difference $[{R}_{n}(\lambda^{\prime\prime})-{R}_{n}(\lambda^{\prime})]$ from Eq. (2) by estimating $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ and $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$ in QMC simulations. Additionally, to avoid overflow/underflow issues, we have to store their logarithms in computers.
@@
The procedure described above is sufficient to derive the relative values of $R_{n}(\lambda)$ at different $\lambda^{\prime\prime}$ when $\lambda^{\prime}$ is fixed at $\lambda_{0}$, called a reference point. This allows for the study of various behaviors, such as locating extreme values and singularities of $R_{n}(\lambda)$. Moreover, if ${R}_{n}(\lambda_{0})$ is known, we can compute the exact values of ${R}_{n}(\lambda)$. Typically, finding such a reference point is straightforward. For example, when $\lambda=\beta\equiv 1/T$, the inverse temperature, we can set $\beta_{0}=0$, which corresponds to a completely disordered system where both classical and quantum correlations are destroyed, giving ${R}_{n}(\beta_{0})=0$. In practice, we use a sufficiently small value (e.g. $\beta_{0}=10^{-8}$) in our simulations, and the adequacy of the choice can be verified by considering a smaller one to check the convergence. Similarly, when $\lambda$ represents the coupling strength of interacting terms in a Hamiltonian related to the entangled boundary, setting $\lambda_{0}=0$ would make the two subsystems completely independent, resulting in ${R}_{n}(\lambda_{0})=0$ again. Therefore, depending on our choice of $\lambda$, both finite and zero temperature properties of $R_{n}$ can be explored. For illustrative purpose, we now set $\lambda=\beta$, and other choices of $\lambda$ can be similarly discussed. By identifying $\lambda_{0}\equiv\beta_{0}\equiv 0$, Eq. (2) can be reformulated as
${R}_{n}(\beta)=\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}-\ln\frac{Z_{n}(0)}{Z_{n}(\beta)}.$ (3)
Importance sampling and annealing scheme.- Although the ideas of reweighting and reference point are simple, simulations become inefficient if $\beta$ is far away from $\beta_{0}$ in the parameter space, due to the significant differences between the distributions $Z_{n}^{T_{B}}(\beta)$ and $Z_{n}^{T_{B}}(\beta_{0})$, requiring an excessive number of samples. To address this, we divide $[\beta_{0},\beta]$ into $m$ small subintervals, i.e.
$\begin{split}\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}=&\sum_{k=1}^{m}\ln\frac{Z_{n}^{T_{B}}(\beta_{k-1})}{Z^{T_{B}}_{n}(\beta_{k})}\\ \ln\frac{Z_{n}(0)}{Z_{n}(\beta)}=&\sum_{k=1}^{m}\ln\frac{Z_{n}(\beta_{k-1})}{Z_{n}(\beta_{k})},\end{split}$ (4)
@@
where $\beta_{m}\equiv\beta$. Then, efficient estimation is ensured when adjacent $\beta_{k}$ values are close to satisfy the importance sampling Neal (2001). In addition, the method’s effectiveness depends on how $[\beta_{0},\beta]$ is divided, as the computational cost also scales with the number of subintervals. Simple approaches, such as equal divisions or geometric progressions with large/small ratios, are inefficient since the required number of subinterval density varies across parameter regions and different system sizes.
@@
To address the problem, we adopt the annealing scheme, which requires a polynomial number of subintervals and has demonstrated excellent performance in practical partition function calculations Ding et al. (2024). The core idea is to approximately restrict $[Z_{n}^{T_{B}}(\beta_{k-1})/Z^{T_{B}}_{n}(\beta_{k})]$ and $[Z_{n}(\beta_{k-1})/Z_{n}(\beta_{k})]$ to be some constant $\epsilon\lesssim 1$ for all sizes and $\beta_{k}$, to ensure the importance sampling. With some calculations, we find that the total number of subintervals $F(L)$ required scales as $F(L)\approx(\beta-\beta_{0})\Lambda nL^{d}/|\ln\epsilon|$, where $d$ is the system dimension, $L$ is the length of system, and $\Lambda$ is a constant related to the energy density. Notably, an incorrect choice of $\Lambda$ just effectively changes the value of $\epsilon$, making $\epsilon$ the sole hyperparameter in the algorithm. Similarly, one can discuss when $\lambda$ is some other parameter in the Hamiltonian, and the division number also scales polynomially with the system size. Further details, including the derivation of $F(L)$ can be found in the SM.
@@
Ground states of tripartite 1D ADHM.-
@@
Figure 2: With $\beta=L$ to extrapolate to the ground state: (a) The variation of $R_{3}(K)$ with respect to the parameter $K$ for the tripartite (1+1)D DHM ring; (b) The central charge $c=1.02(2)$ of the underlying CFT for the Heisenberg model is fitted; (c) The variation of $R_{3}(h)$ with respect to the parameter $h$ for the (1+1)D TFIM under half bipartition; (d) The central charge of the underlying CFT at the critical point is fitted to be $c=0.51(3)$. Notably, the data of $L=16$, which exhibits significant finite-size effects and clearly deviates from the trend, is excluded from the fitting.
@@
for a (1+1)D CFT, where $l$ is the length of subsystems $A$ and $B$. With Eq. (6), the central charge is obtained to be $c=0.51(3)$, which also accords with the theoretical value. Physically, for both the tripartite 1D ADHM above and this bipartite 1D TFIM, the logarithmic correction of $R_{n}$ at the CFT points, which is beyond the area law, reflects the existence of long-range entanglement.
Figure 3: For the 2D TFIM on the square lattice: (a) fixing $h=3.04438$, the variation of $R_{3}(J)$ as a function of $J$, where $\beta=L$ to extrapolate to the groud state; (b) the area law correction $-\gamma(J)$ of $R_{3}(J)$ for $J\in[0,1.2]$, and the inset shows the fitting of the area law at $J=J_{c}$, where $-\gamma(J_{c})=0.02(5)$; (c) fixing $J=1$ and $h=2.75$, the variation of $R_{3}(\beta)$ as a function of $\beta$. (d) the area law correction $-\gamma(\beta)$ of $R_{3}(\beta)$ for $\beta\in(0,0.4]$, and the inset shows the fitting of the area law at $\beta=\beta_{c}/3$, where $-\gamma(\beta_{c})=0.002(7)$;
@@
For the ground states, Fig. 3(a) illustrates the behavior of $R_{3}(J)$ as a function of $J$ for different system sizes. Similar to the 1D TFIM, $R_{3}(J)$ reaches a maximum and diverges at the QCP, indicating that the QCP exhibits stronger quantum correlations compared to the gapped phases. In the gapped phases, no long-range entanglement is expected, and we should have $R_{3}\sim aL-\gamma$ with $\gamma=0$. Near the critical point, we observe that the area law breaks down, with $R_{3}$ for small system sizes exceeding that of larger sizes. However, as the system moves further from the criticak point, the area law is restored. If we forcibly fit $R_{3}$ to the area law, the constant term $\gamma$ fluctuates significantly near the critical point, as shown in Fig. 3(b). In other regions, it remains zero within error bars. In principle, $R_{n}$ can be understood as the difference of free energies between two space-time manifolds with different boundary conditions. Since the only distinction lies at the boundary, it should follow an area law. The deviation from the area law in our numerical data is due to finite-size effects in both space and time, as the gaplessness related to the phase transition requires larger system sizes. Further details on finite-size effects are provided in the SM. Notably, RN inherently removes the contribution of ground state degeneracy due to the denominator in its definition, whereas EE captures it in the correction term as $\gamma=-\ln(N_{\text{deg}})$, where $N_{\text{deg}}$ denotes the ground state degeneracy Stéphan et al. (2009). In this example of ferromagnetic phase, we have $\gamma=0$.
@@
For the Gibbs states, we take $\beta_{0}=0$ to be the reference point, and Fig. 3(c) shows that $R_{3}(\beta)$ increases monotonically from the infinite high temperature $\beta_{0}$ to $\beta=0.4$. Notably, the three replicas effectively scale the inverse temperature by a factor of $1/3$ in partition function $Z_{3}$, and the related critical properties are shifted to $\beta_{c}/3$  [^1] The term $\mathrm{tr}(\rho^{3})\propto\text{tr}(e^{-3\beta H})$ will lead a singularity at $\beta_{c}/3$. This has been confirmed in the calculation of the divergence of the first derivative of $a$, the area-law coefficient of $R_{3}$, in the quantum spherical model and 2D TFIM Wu et al. (2020a). As seen in Fig. 3(c), at the critical point $\beta_{c}/3$, no significant behavior associated with the TCP is observed for $R_{3}$ in our calculations. Furthermore, we find that $R_{3}$ perfectly obeys the area law and $\gamma\sim 0$ for all the $\beta$ we computed, as shown in Fig. 3(d). This observation is consistent with the fact that the phase transition is driven by thermal fluctuations, while quantum phase transitions in the 1D TFIM, whose QCP shares the universality class with the TCP, is governed by quantum fluctuations. Consequently, no quantum effects can contribute to the generation of a nonzero and nonlocal term $\gamma$ in $R_{3}$ Lu et al. (2020); Lu and Grover (2020b); Lu (2024). Although this conclusion may seem straightforward, it was previously difficult to observe. Moreover, the increase in $R_{3}$ when decreasing the temperature reflects an enhancement of quantum properties. However, the quantum correlations remain short range ($\gamma\sim 0$), indicating that the phases beyond or below the TCP are both trivial in terms of quantum correlation.
@@
In addition, we discuss the derivatives of $R_{n}$ with respect to some parameter $\lambda$. Given that $R_{n}(\lambda)=-\ln Z_{n}^{T_{B}}(\lambda)+\ln Z_{n}(\lambda)$, it follows that if the $k$th-order derivatives of both effective free energy terms are nonsingular at $\lambda$, then $d^{k}R_{n}(\lambda)/d\lambda^{k}$ must also remain nonsingular. Morover, since the difference between the two systems associated with $Z_{n}^{T_{B}}$ and $Z_{n}$ arises solely from their boundary conditions, their critical points $\lambda_{c}$ should remain identical in the thermodynamic limit. Consequently, any singularity in a higher-order derivative of $R_{n}$ must occur precisely at $\lambda_{c}$.
@@
Summary and discussions.- We have introduced an efficient algorithm for calculating Rényi negativity in quantum Monte Carlo simulations, built on the reweight-annealing framework. This algorithm is easy to implement, has polynomial computational complexity, and is naturally parallelizable. By enabling the calculation of Rényi negativity along a full parameter trajectory in parameter space, our method proves highly versatile. It is not only valuable for investigating entanglement behavior across different phases and near critical points, but also for fitting correction terms in area laws to extract universal quantities such as central charge and topological order in quantum many-body systems. We have validated the approach in various scenarios, including 1D and 2D systems, ground and Gibbs states, as well as bipartition and tripartition setups. Our results of the Rényi negativity further reveals a detailed understanding of the intrinsic mechanisms governing thermal and quantum critical points in many-body systems.

### v3

@@ Abstract
Introduction.- The increasing connection between condensed matter physics and quantum information science have significantly deepened our understanding of quantum many-body behaviors. Beyond the conventional Landau-Ginzburg-Wilson framework of symmetry breaking and phase transitions Sachdev (1999); Lifshitz and Pitaevskii (2013), quantum entanglement and other intricate microcosmic degrees of freedom have become central in explaining various exotic emergent phenomena Amico et al. (2008); Laflorencie (2016); Zeng et al. (2019). In a bipartite composite system, the entanglement of a pure ground state can be quantified by entanglement entropy (EE), which captures universal information of the system including critical behaviors Kallin et al. (2014); Helmes and Wessel (2014); Zhao et al. (2022a, b), continuous symmetry breaking Metlitski and Grover (2011); D’Emidio (2020); Deng et al. (2023); Wang et al. (2024a), the underlying conformal field theory (CFT) Calabrese and Lefevre (2008); Fradkin and Moore (2006); Casini and Huerta (2007); Ji and Wen (2019); Tang and Zhu (2020) and topological order Kitaev and Preskill (2006); Levin and Wen (2006); Isakov et al. (2011); Nussinov and Ortiz (2009a, b). However, EE fails as a measure of entanglement in mixed states, which frequently arise in many-body physics, such as in finite-temperature Gibbs states, systems with disconnected partitions, and open quantum systems Lifshitz and Pitaevskii (2013); Laflorencie (2016); Yan et al. (2018).
@@
In recent years, some numerical approaches have been developed to calculate RN or the moments Alba (2013); Shim et al. (2018); Chung et al. (2014a); Wu et al. (2020a); Wang and Xu (2025). However, particularly for large-scale interacting spin/boson systems, there are few quantum Monte Carlo (QMC) works Chung et al. (2014b); Wu et al. (2020b), mainly because the incremental trick commonly used in calculating Rényi EE Hastings et al. (2010); Humeniuk and Roscilde (2012); D’Emidio (2020) becomes much more complex and difficult in the case of RN, since the lowest order of nontrivial RN is third and the boundary condition of imaginary time needs to be incrementally changed from a glued space-time manifold to a twistedly glued one. Therefore, developing an efficient QMC algorithm that is both easy to implement and capable of accurately computing RN is a crucial yet challenging task.
Furthermore, the absence of such algorithms has left many important questions about mixed-state entanglement unexplored. For example, one might naturally assume that the infinite quantum correlation length at a quantum critical point necessarily implies long-range entanglement. However, this is not always the case. Quantum correlation length does not directly measure entanglement: for instance, gapped topological phases can exhibit long-range entanglement despite having finite correlation lengths. Meanwhile, the classical Ising model has no entanglement even at the thermal critical point with infinite correlation length. Nevertheless, the thermal phase transition of a quantum system has both thermal and quantum fluctuations. Due to the existence of quantum fluctuation, the $T_{c}$ of thermal phase transition of a quantum model is usually lower than a pure classical model. It can thus be seen that quantum fluctuations are involved in the thermal phase transition of quantum systems. A fundamental question is what role the entanglement plays in this case.
In this letter, we have filled the technical gap by presenting an efficient QMC algorithm within the reweight-annealing frame Ding et al. (2024); Wang et al. (2024b); Jiang et al. (2024); Ding et al. (2025) and explored several fundamental questions about the entanglement in mixed states.
@@
Figure 1: The horizontal axis represents spatial spins (depicted as circles), while the vertical axis corresponds to imaginary time. Spins of the same color are connected (or glued) along the imaginary time direction by straight lines of the same color. (a) Manifold for $Z_{3}$: The three replicas are sequentially connected along the imaginary time direction for the system $A\cup B$. The environment, denoted by $\overline{A\cup B}$, is traced out by imposing periodic boundary conditions on each replica. (b) manifold for $Z_{3}^{T_{B}}$: The connections among three replicas in subsystem $A$ (the blue part) and the environment $\overline{A\cup B}$ (the yellow part) remain identical to that in the left manifold. However, in subsystem $B$, the connections are not in sequence due to the operation $T_{B}$.
@@
The ratio $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ in Eq. (2) can be further interpreted as a reweighting operation Ferrenberg and Swendsen (1988, 1989); Troyer et al. (2004); Dai and Xu (2025). If express $Z_{n}^{T_{B}}(\lambda)=\sum_{i}W_{i}(\lambda)$, where $\{W_{i}(\lambda)\}$ are the weights, then the ratio can be viewed as the average of weight ratios, since $Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})=\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$, where $W(\lambda)\in\{W_{i}(\lambda)\}$ and the expectation is computed with respect to samples from $Z_{n}^{T_{B}}(\lambda^{\prime\prime})$. The specific form of the estimator $\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$ depends on the QMC method we use, which transforms the quantum degrees of freedom in $Z_{n}^{T_{B}}(\lambda^{\prime})$ to some classical ones. For illustration, we use the stochastic series expansion (SSE) method in this letter Sandvik (1998, 2003); Yan et al. (2019); Yan (2022), and the corresponding estimators can be found in the Supplemental Material sup (see also references Sandvik (1998, 2003); Yan et al. (2019); Yan (2022); Zhou et al. (2022, 2023); Yan et al. (2023); Calabrese et al. (2013a); Laflorencie (2016) therein). Similarly, we can apply the same reasoning to $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$. It is important to emphasize that the view of reweighting is crucial, as it allows us to estimate the difference $[{R}_{n}(\lambda^{\prime\prime})-{R}_{n}(\lambda^{\prime})]$ from Eq. (2) by estimating $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ and $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$ in QMC simulations. Additionally, to avoid overflow/underflow issues, we have to store their logarithms in computers.
@@
The procedure described above is sufficient to derive the relative values of $R_{n}(\lambda)$ at different $\lambda^{\prime\prime}$ when $\lambda^{\prime}$ is fixed at $\lambda_{0}$, called a reference point. This allows for the study of various behaviors, such as locating extreme values and singularities of $R_{n}(\lambda)$. Moreover, if ${R}_{n}(\lambda_{0})$ is known, we can compute the exact values of ${R}_{n}(\lambda)$. Typically, finding such a reference point is straightforward. For example, when $\lambda=\beta\equiv 1/T$, the inverse temperature, we can set $\beta_{0}=0$, which corresponds to a completely disordered system where both classical and quantum correlations are destroyed, giving ${R}_{n}(\beta_{0})=0$. In practice, we use a sufficiently small value (e.g. $\beta_{0}=10^{-8}$) in our simulations, and the adequacy of the choice can be verified by considering a smaller one to check the convergence. Similarly, when $\lambda$ represents the coupling strength of interacting terms in a Hamiltonian related to the entangled boundary, setting $\lambda_{0}=0$ would make the two subsystems completely independent, resulting in ${R}_{n}(\lambda_{0})=0$ again. Therefore, depending on our choice of $\lambda$, both finite and zero temperature properties of $R_{n}$ can be explored. For illustrative purpose, we now set $\lambda=\beta$, and other choices of $\lambda$ can be similarly discussed. By identifying $\lambda_{0}\equiv\beta_{0}\equiv 0$, Eq. (2) can be reformulated as ${R}_{n}(\beta)=\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}-\ln\frac{Z_{n}(0)}{Z_{n}(\beta)}$.
@@
Importance sampling and annealing scheme.- Although the ideas of reweighting and reference point are simple, simulations become inefficient if $\beta$ is far away from $\beta_{0}$ in the parameter space, due to the significant differences between the distributions $Z_{n}^{T_{B}}(\beta)$ and $Z_{n}^{T_{B}}(\beta_{0})$, requiring an excessive number of samples. To address this, we divide $[\beta_{0},\beta]$ into $m$ small subintervals, i.e. $\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}=\sum_{k=1}^{m}\ln\frac{Z_{n}^{T_{B}}(\beta_{k-1})}{Z^{T_{B}}_{n}(\beta_{k})}$ and $\ln\frac{Z_{n}(0)}{Z_{n}(\beta)}=\sum_{k=1}^{m}\ln\frac{Z_{n}(\beta_{k-1})}{Z_{n}(\beta_{k})}$, where $\beta_{m}\equiv\beta$. Then, efficient estimation is ensured when adjacent $\beta_{k}$ values are close to satisfy the importance sampling Neal (2001). In addition, the method’s effectiveness depends on how $[\beta_{0},\beta]$ is divided, as the computational cost also scales with the number of subintervals. Simple approaches, such as equal divisions or geometric progressions with large/small ratios, are inefficient since the required number of subinterval density varies across parameter regions and different system sizes.
@@
To address the problem, we adopt the annealing scheme, which requires a polynomial number of subintervals and has demonstrated excellent performance in practical partition function calculations Ding et al. (2024). The core idea is to approximately restrict $[Z_{n}^{T_{B}}(\beta_{k-1})/Z^{T_{B}}_{n}(\beta_{k})]$ and $[Z_{n}(\beta_{k-1})/Z_{n}(\beta_{k})]$ to be some constant $\epsilon\lesssim 1$ for all sizes and $\beta_{k}$, to ensure the importance sampling. With some calculations, we find that the total number of subintervals $F(L)$ required scales as $F(L)\approx(\beta-\beta_{0})\Lambda nL^{d}/|\ln\epsilon|$, where $d$ is the system dimension, $L$ is the length of system, and $\Lambda$ is a constant related to the energy density. Notably, an incorrect choice of $\Lambda$ just effectively changes the value of $\epsilon$, making $\epsilon$ the sole hyperparameter in the algorithm. Similarly, one can discuss when $\lambda$ is some other parameter in the Hamiltonian, and the division number also scales polynomially with the system size. Further details, including the derivation of $F(L)$ can be found in the Supplemental Material sup .
@@
Ground states of tripartite 1D antiferromagnetic dimerzied Heisenberg model (ADHM).-
@@
Figure 2: With $\beta=L$ to extrapolate to the ground state: (a) The variation of $R_{3}(K)$ with respect to the parameter $K$ for the tripartite (1+1)D ADHM ring; (b) The central charge $c=1.02(2)$ of the underlying CFT for the Heisenberg model is fitted; (c) The variation of $R_{3}(h)$ with respect to the parameter $h$ for the (1+1)D transverse field Ising model (TFIM) under half bipartition; (d) The central charge of the underlying CFT at the critical point is fitted to be $c=0.51(3)$. Notably, the data of $L=16$, which exhibits significant finite-size effects and clearly deviates from the trend, is excluded from the fitting.
@@
for a (1+1)D CFT, where $l$ is the length of subsystems $A$ and $B$. With Eq. (4), the central charge is obtained to be $c=0.51(3)$, which also accords with the theoretical value. In both the tripartite 1D ADHM and the bipartite 1D TFIM, the logarithmic corrections to $R_{n}$ at the CFT points deviate from the expected relation $R(2l)=2R(l)$, signaling the presence of long-range entanglement Lu et al. (2020); Lu and Grover (2020b); Lu (2024). In contrast, a system with only short-range entanglement would yield an extensive $R_{n}$, with no subleading terms beyond the area law.
Figure 3: For the 2D TFIM on the square lattice: (a) fixing $h=3.04438$, the variation of $R_{3}(J)$ as a function of $J$, where $\beta=L$ to extrapolate to the ground state; (b) the area law correction $-\gamma(J)$ of $R_{3}(J)$ for $J\in[0,1.2]$. The inset panel shows the fitting of area law with the maximum value of $R_{3}(L,J)$ for each $L$, and the area law coefficient is $\gamma=-0.225(49)$; (c) fixing $J=1$ and $h=2.75$, the variation of $R_{3}(\beta)$ as a function of $\beta$. (d) the area law correction $-\gamma(\beta)$ of $R_{3}(\beta)$ for $\beta\in(0,0.4]$, and the inset shows the fitting of the area law at $\beta=\beta_{c}/3$, where $-\gamma(\beta_{c})=0.002(7)$;
@@
For the ground states, Fig. 3(a) illustrates the behavior of $R_{3}(J)$ as a function of $J$ for different system sizes. Similar to the 1D TFIM, $R_{3}(J)$ reaches a maximum and diverges at the QCP, indicating that the QCP exhibits stronger quantum correlations compared to the gapped phases. In the gapped phases, no long-range entanglement is expected, and we should have $R_{3}\sim aL-\gamma$ with $\gamma=0$. Near the critical point, we observe that the area law breaks down, with $R_{3}$ for small system sizes exceeding that of larger sizes, as shown in Fig. 3(a). However, as the system moves further from the critical point, the area law is restored. If we forcibly fit $R_{3}$ to the area law, the constant term $\gamma$ fluctuates significantly near the critical point, as shown in Fig. 3(b). In other regions, it remains zero within error bars. In principle, $R_{n}$ can be understood as the difference of free energies between two space-time manifolds with different boundary conditions. Since the only distinction lies at the boundary, it should follow an area law. The deviation from the area law in our numerical data is due to finite-size effects in both space and time, as the gaplessness related to the phase transition requires larger system sizes. Further details on finite-size effects are provided in the Supplemental Material sup .
To reduce finite-size effects, we use the peaks of $R_{3}(J)$ across different system sizes to fit the area law at the QCP, as shown in the inset of Fig. 3(b) [^1] For large sizes, this result converges to that obtained by fitting at a fixed parameter point.. Its constant correction extracted from the fit is $\gamma=-0.225(49)$  [^2] Notice we use $R_{3}\sim aL-\gamma$ to fit, thus the final constant correction implies a positive sub-leading correction., which signifies a breakdown of the local behavior of RN and reflects the long-range entanglement at the QCP here. This constant correction is consistent with the theoretical prediction Metlitski et al. (2009). In addition, RN inherently removes the contribution of ground state degeneracy due to the denominator in its definition, whereas EE captures it in the correction term as $\gamma=-\ln(N_{\text{deg}})$, where $N_{\text{deg}}$ denotes the ground state degeneracy Stéphan et al. (2009). In this example of ferromagnetic phase, we have $\gamma=0$ while $N_{\text{deg}}=2$. This reflects the advantage of RN in computing mixed-state properties, where it naturally filters out degeneracy effects.
@@
For the Gibbs states, Fig. 3(c) shows that $R_{3}(\beta)$ increases monotonically from the infinite high temperature $\beta_{0}$ to $\beta=0.4$. Notably, the three replicas effectively scale the inverse temperature by a factor of $1/3$ in partition function $Z_{3}$, and the related critical properties are shifted to $\beta_{c}/3$  [^3] The term $\mathrm{tr}(\rho^{3})\propto\text{tr}(e^{-3\beta H})$ will lead a singularity at $\beta_{c}/3$. This has been confirmed in the calculation of the divergence of the first derivative of $a$, the area-law coefficient of $R_{3}$, in the quantum spherical model and 2D TFIM Wu et al. (2020a). As seen in Fig. 3(c), at the critical point $\beta_{c}/3$, no significant behavior associated with the TCP is observed for $R_{3}$ in our calculations. Furthermore, we find that $R_{3}$ perfectly obeys the area law and $\gamma\sim 0$ for all the $\beta$ we computed, as shown in Fig. 3(d). Despite choosing $h=2.75$, close to the QCP and dominated by strong quantum fluctuations, our findings indicate that thermal phase transitions remain short-range entangled, highlighting a key distinction from the long-range entanglement inherent in quantum criticality. Moreover, the increase in $R_{3}$ when decreasing the temperature reflects an enhancement of quantum properties. However, the quantum correlations remain short range ($\gamma\sim 0$), indicating that the phases beyond or below the TCP are both trivial in terms of quantum correlation.
@@
In addition, we discuss the derivatives of $R_{n}$ with respect to some parameter $\lambda$. Given that $R_{n}(\lambda)=-\ln Z_{n}^{T_{B}}(\lambda)+\ln Z_{n}(\lambda)$, it follows that if the $k$th-order derivatives of both effective free energy terms are non-singular at $\lambda$, then $d^{k}R_{n}(\lambda)/d\lambda^{k}$ must also remain non-singular. Moreover, since the difference between the two systems associated with $Z_{n}^{T_{B}}$ and $Z_{n}$ arises solely from their boundary conditions, their critical points $\lambda_{c}$ should remain identical in the thermodynamic limit. Consequently, any singularity in a higher-order derivative of $R_{n}$ must occur precisely at $\lambda_{c}$.
@@
Summary and discussions.- We have introduced an efficient algorithm for calculating Rényi negativity in quantum Monte Carlo simulations, built on the reweight-annealing framework. This algorithm is easy to implement, has polynomial computational complexity, and is naturally parallelizable. By enabling the calculation of Rényi negativity along a full parameter trajectory in parameter space, our method proves highly versatile. It is not only valuable for investigating entanglement behavior across different phases and near critical points, but also for fitting correction terms in area laws to extract universal quantities such as central charge and topological order in quantum many-body systems. We have validated the approach in various scenarios, including 1D and 2D systems, ground and Gibbs states, as well as bipartition and tripartition setups. Our results of the Rényi negativity further reveals a detailed understanding of the intrinsic mechanisms governing thermal and quantum critical points in many-body systems. An intriguing direction for future research is to investigate the role of entanglement in strong-to-weak spontaneous symmetry breaking in mixed quantum states Lessa et al. (2025a, b).